OpenStudy (anonymous):

x = 0.8x + 72 I dont know what to do with the x thats next 0.8

OpenStudy (anonymous):

x = 0.8x + 72 - 0.8 x on both sides 0.2x = 72 divide 0.2 both sides ..... profit

OpenStudy (anonymous):

So the x just goes away in a sense?

OpenStudy (anonymous):

when you divide it?

OpenStudy (anonymous):

1 x = 0.8x + 72 - 0.8x -0.8x ----------------------- 0.2x = 0 + 72 oh 0.2x = 72 divide both sides by 0.2 x = 360 you can't do anything to x, so it stays there.

OpenStudy (anonymous):

OKay thanks for the help :)